OpenStudy (anonymous):

When writing of the sqrt of -13 is the negative sign placed under the radicand or outside of it?

myininaya (myininaya):

\[\sqrt{-1}=i\]

OpenStudy (anonymous):

13i

OpenStudy (dumbcow):

neither, replace the neg sign with an "i" on outside :)

myininaya (myininaya):

\[\sqrt{-a} =i \sqrt{a} , a \ge 0\]

OpenStudy (anonymous):

I am not doing complex numbers yet... so I don't think I can use i

OpenStudy (dumbcow):

in that case...its No real solution and it doesn't matter, leave the neg inside. sqrt(-13)

OpenStudy (dumbcow):

\[\sqrt{-13} \ne -\sqrt{13}\]

OpenStudy (anonymous):

am solving x^2+10x+38 =0... so I say it has no real soln's?

OpenStudy (dumbcow):

correct

myininaya (myininaya):

no real solution if b^2-4ac<0

OpenStudy (anonymous):

at what point do I say that at the end or as soon as I come across sqrt{-13}

OpenStudy (dumbcow):

as soon as you come across it...whenever discriminant is negative like myininaya said

OpenStudy (anonymous):

okay, thank you all so much!!